A new sports car model has defective brakes 15 percent of the time and a defective steering mechanism 5 percent of the time. Let’s assume (and hope) that these problems occur independently. If one or the other of these problems is present, the car is called a “lemon.” If both of these problems are present, the car is a “hazard.” Your instructor purchased one of these cars yesterday.
What is the probability it is:
a) hazard ?
b) a lemon?

Answer by ikleyn(52781)

.

P(hasard) = 0.15*0.05 = 0.0075.                  ANSWER


P(lemon) = 0.15 + 0.05 - 0.15*0.05 = 0.1925.     ANSWER